OpenStudy (anonymous):

Find the Limit:

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty}(\frac{ 5x-3 }{ 5x+7 })^{3x+1}\]

OpenStudy (anonymous):

take the log take the limit exponentiate

OpenStudy (anonymous):

I am allowed to use L'Hospital's rule. I am not sure how I would use it though...

OpenStudy (anonymous):

So I get it's (infinity/infinity)^(infinity) but I am not sure how to go from there...

OpenStudy (anonymous):

Yep... I am not sure how that would help though...

OpenStudy (anonymous):

\[(3x+1)\ln(\frac{5x-3}{5x+7})\] now you \(\infty\times 0\)

OpenStudy (anonymous):

Yep... I got there.

OpenStudy (anonymous):

rewrite as \[\frac{\ln(\frac{5x-3}{5x+7})}{\frac{1}{3x+1}}\] now you have \[\frac{0}{0}\] and now you can use l'hopital

OpenStudy (anonymous):

Hmm Okay... Let's see...

OpenStudy (anonymous):

probably easier if you write \[\frac{\ln(5x-3)-\ln(5x+7)}{\frac{1}{3x+1}}\]

OpenStudy (anonymous):

Okay Thanks!

OpenStudy (anonymous):

Yep. I know.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

@satellite73 : Okay so I have: \[\lim_{x \rightarrow \infty} \frac{ \frac{ 5 }{ 5x-3 }-\frac{ 5 }{ 5x+7 } }{ \frac{ -2 }{ (3x+1)^2 } }\]

OpenStudy (anonymous):

denominator is wrong

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} \frac{ \frac{ 5 }{ 5x-3 }-\frac{ 5 }{ 5x+7 } }{ \frac{ -3 }{ (3x+1)^2 } }\]

OpenStudy (anonymous):

Chain rule. Right.

OpenStudy (anonymous):

ugly retricealgebra from here on in have fun, but you will get a rational expression so limit is same as horizontal asymptote

OpenStudy (anonymous):

:( . Fair enough....

OpenStudy (anonymous):

actually it is not that bad

OpenStudy (anonymous):

Yep. I think I see why.

OpenStudy (anonymous):

\[-\frac{(3x+1)^2}{3}\times \frac{50}{(5x-3)(5x+7)}\] and not need to multiply out, just look at the ratio of the leading coefficients

OpenStudy (anonymous):

i mean you have to be a bit careful but the leading coefficient of the numerator is \(9\times 50\) and the leading coefficient of the denominator is \(-3\times 25\) so you the ratio is \(-6\)

OpenStudy (anonymous):

and don't forget to exponentiate at the end

OpenStudy (anonymous):

Hmm... Okay.

OpenStudy (anonymous):

So I got 1/e^6 or e^-6.

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

:) . Thanks a million :) .